A typical jet liner lands at a speed of 100m/s. while braking, it has an acceleration of -5.2m/s squared. (a) How long does it take to stop? (b) What is the minimum length of the landing strip under these conditions?


1
Expert's answer
2021-04-19T17:06:48-0400

(a) Find the time:


"t=\frac{v_f-v_i}{a}=\frac{0-100}{-5.2}=19.2\text{ s}."

(b) The distance is


"d=\frac{v_f^2-v_i^2}{2a}=\frac{0^2-100^2}{2·(-5.2)}=961.5\text{ m}."